OpenStudy (anonymous):

10 Find the distance of point P (-7,0,5) from the straight line passing through points A(-2,0,1) and B(3,2,2) .

OpenStudy (amistre64):

slope between A and B maybe, then perp it ... might be easier said then done in that way tho

OpenStudy (amistre64):

-2 0 1 -3-2-2 ------- -5 -2 -1 .... <5,2,1> is the vector between our points

OpenStudy (amistre64):

line: x=-2+5t y= 0+2t z= 1+ t plane with out given point is: 5(x--7)+2(y-0)+1(z-5)=0 this line and this plane intersect at: 5(-2+5t+7)+2(0+2t-0)+1(1+ t-5)=0 25+25t+4t-4+ t=0 (25-4)+t(25+4+1)=0 (21)+t(30)=0 ; t=7/10 maybe x=-2+5.7/10= 15/10 y= 0+2.7/10= 14/10 z= 1+ 7/10 = 17/10 so the distance between this point and the other might be it if I did the mathing roght

OpenStudy (amistre64):



OpenStudy (anonymous):

I,m still puzzled, whats the distance. the value for t

OpenStudy (amistre64):

the distance in my case would be the distance formula using the point given and the found point; t was just used to find the needed point

OpenStudy (amistre64):

a simpler method might be to employ a sin of the angle between vectors tho

OpenStudy (amistre64):

AB and AP are 2 vectors; find the angle between them and the distance is AP sin(a)

OpenStudy (anonymous):

Thats how they do it but i dont really know how do u.

OpenStudy (amistre64):

I do it in whatever manner makes sense to me .... seeing how I do not have any material to base a solution on I have to pretty much make one up that will fit the bill

OpenStudy (anonymous):

i try to understand how u did it. but whats the distance that u got

OpenStudy (amistre64):

if my math is right, we can take our discovered point and subtract it from our known point -7 , 0 , 5 -15/10, -14/10, -17/10 -70/10, 0/10, 50/10 -15/10, -14/10, -17/10 ----------------------- -85/10, -14/10, 33/10 square add amd sqrt the results sqrt(85^2+14^2+33^2) /10 would be our distance, about 9.22 we can check this with the sine method to be sure :)

OpenStudy (amistre64):

P (-7,0,5) B(3,2,2) . -A(-2,0,1) -A(-2,0,1) -------------------- a<-5,0,4> b<5,2,1> |a| = sqrt(41) |b|=sqrt(30) a<-5,0,4> .b<5,2,1> ----------- -25+0+4 = -21 cos(t) = -21/sqrt(41*30) t = arccos(-21/sqrt(41*30)) |a| sin(t) = our distance than d = sqrt(41) sin(arccos(-21/sqrt(41*30))) = 5.13 so that means my math is off someplace, or ive forgoteen a step :/

OpenStudy (amistre64):